Which of the following is an ordered pair solution of y=-1/2x+1/4?
A. (2,-1/2)
B. (1,1/4)
C. (-2,5/4)
D. (-1,-5/4)

Answer by checkley77(12844)

ANS. C)
y=-1/2x+1/4
PROOF:
5/4=-1/2*-2+1/4
5/4=2/2+1/4
5/4=4/4+1/4
5/4=5/4